1) Through how many degrees does the minute hand of a clock turn during:
a)60 minutes b)40 minutes c)35 minutes
2) Through how many degrees does the hour hand of a clock turn during:
a)1 hour b)0.5 hour c)15/2 hour d)7/2 hour

Answer by sachi(548)

1) the min hand when turns 1 complete circle on the clock it covers 360 degree in 60 mins
a) so in 60 mins the min hand moves 360 degree
b) in 40 mins (360/60)40=240 degrees
c) in 35 mins (360/60)35=210 degrees
2) the hour hand when turns 1 complete circle on the clock it covers 360 degree in 12 hrs
a) so in 1 hr the hr hand moves 360/12=30 degree
b) in 0.5 hr (360/12)*0.5=15 degrees
c) in 15/2 hr (360/12)*15/2=225 degrees
d) in 7/2 hrs the hr hand moves (360/12)*7/2=105 degree
ans